Compute the real interest rate in each case:

(a) i=4%, π e = 2%

(b) i=16%, π e = 13%

(c) i=2%, π e = 0%
Answer #1

Solution:

The required formula is: real interest rate = nominal interest rate - expected inflation rate

r = i - pie

So, we get

(a) i = 4%, pie = 2%

r = 4 - 2 = 2%

(b) i = 16%, pie = 13%

r = 16 - 13 = 3%

(c) i = 2%, pie = 0%

r = 2 - 0 = 2%